# A Generalized Label Shift Perspective for Cross-Domain Gaze Estimation

### v1

@@ Abstract
Aiming to generalize the well-trained gaze estimation model to new target domains, Cross-domain Gaze Estimation (CDGE) is developed for real-world application scenarios. Existing CDGE methods typically extract the domain-invariant features to mitigate domain shift in feature space, which is proved insufficient by Generalized Label Shift (GLS) theory. In this paper, we introduce a novel GLS perspective to CDGE and modelize the cross-domain problem by label and conditional shift problem. A GLS correction framework is presented and a feasible realization is proposed, in which a importance reweighting strategy based on truncated Gaussian distribution is introduced to overcome the continuity challenges in label shift correction. To embed the reweighted source distribution to conditional invariant learning, we further derive a probability-aware estimation of conditional operator discrepancy. Extensive experiments on standard CDGE tasks with different backbone models validate the superior generalization capability across domain and applicability on various models of proposed method.
@@ 1 Introduction
Gaze estimation (GE) is crutial for understanding human attention in many areas, such as human-robot interaction yang2023natural , virtual reality wagner2024gaze and medical analysis ma2024eye . Recently, leveraging the capabilities of deep models, appereance-based gaze estimation zhang2015appearance ; nagpure2023searching ; liu2024test has attracted wide attention due to its low device requirements and end-to-end workflows. However, the great performance of deep models relies on the identical assumption of training and testing data distributions, which is hard to fulfill in real-world applications. When being applied across data domain, models’ performance usually degrade dramatically due to the domain shift caused by the changes of subjects, environment, etc. Therefore, dealing with the cross-domain gaze estimation (CDGE) problem is crutial for expanding application prospects of gaze estimation.
@@
Figure 1: Illustration of GLS correction framework. The DG and UDA methods typically learn the invariant representation across domains. Differently, the proposed GLS correction framework consider both the label shift correction and the probability-aware invariant representation learning.
@@ 2 Preliminary
Problem Setup. Let $X$ and $Y$ be image and gaze variables defined on $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{D}^{s}=\{(x^{s},y^{s})\}$ and $\mathcal{D}^{t}=\{(x^{t},y^{t})\}$ be source and target domains with elements sampled from variables following different distributions, i.e., $(x^{s},y^{s})\sim(X^{s},Y^{s})\sim P^{s}_{XY}$ and $(x^{t},y^{t})\sim(X^{t},Y^{t})\sim P^{t}_{XY}$. For distribution $P$, the lowercase $p$ is the probability density function. Denote $g:X\to Z\in\mathcal{Z}$ as the feature transformation, $h:Z\to Y$ as the gaze predictor, then a learning model can be regarded as a tuple $(g,h)$. Let $l(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a loss function, then the CDGE problems aim to learn the model that minimizes the prediction error in target domain as
@@
Existing CDGE methods can be roughly categorized as DG and UDA methods. (1) DG methods. Puregaze cheng2022puregaze and GazeCon xu2023learning tackle the CDGE problems by gaze feature purification bao2024unsupervised , which eliminate the gaze-irrelevant factors to obtain gaze-specific feature. GFAL xu2024gaze utilizes the gaze frontalization process and CUDA-GHR jindal2023cuda adopts gaze redirection yu2020unsupervised as auxiliary learning task to improve the generalization capability of features. An attractive advantage of DG methods is that they don’t need target domain samples and the gaze features can be generalized to unseen target domain. However, the improvement is usually limited due to the absence of knowledge from target domain. (2) UDA methods. Given some unlabeled samples from target domain, the UDA methods enhance the model performance in target domain by distribution alignment. DAGEN guo2020domain , PnP-GA liu2021generalizing and PnP-GA+ liu2024pnp align the source and target distributions by different metrics while CRGA wang2022contrastive and HFC liu2024gaze minimize the discrepancy in a contrastive learning paradigm. Recently, UnReGA cai2023source , DUCA zhang2024domain propose to extract domain-invariant representations by reducing the uncertainty within features. Although these DG and UDA methods differ in technology, they essentially focus on domain-invariant representation learning.
GLS theory. A joint distribution $P_{XY}$ can be factorized as $P_{XY}=P_{X|Y}P_{Y}$. Then the discrepancy between $P^{s}_{XY}$ and $P^{t}_{XY}$ is naturally deduced to the following two shifts:
[…]
@@ 3.1 A GLS Perspective of CDGE problem
Figure 2: A GLS perspective of CDGE problem. (a) The label probability functions differ in functions support (colored area) and probability values (color degree), resulting the label shift between domains. (b) Conditional distributions of the same gaze label may differ due to the difference of identity, background and illuminations, etc.
@@
Conditional shift. The difference in sample level is more intuitive. In different scenarios the appereance image may differ significantly due to the background, illuminations, etc, leading to the discrepancy of conditional distributions, i.e.,
@@ 3.2 GLS correction framework for CDGE
  • Estimating the importance weight function $\omega(y)$ by
[…]
    such that $P^{t}_{Y}=\omega(y)P^{s}_{Y}\triangleq P^{s}_{Y^{\omega}}$.
@@
[…]
    to align the feature distributions $P^{s}_{Z|Y^{\omega}}$ and $P^{t}_{Z|Y}$.
@@
The terms {$\mathcal{L}_{\{\}}(\cdot,\cdot)$ denotes the learning Objectives in each step. By considering different steps and distance on distributions, the framework are connected with existing CDGE methods.
@@
Figure 3: Overview of the proposed GLS correction method. (a) Reweighting source label distribution with bivariate Gaussian distribution estimated by pseudo target labels. The reweighted label distribution is utilized in both conditional alignment and task-specific training. (b) Conditional invariant representation learning. Two sets of conditional distribution are map to RKHS and the discrepancy is measured by the PCOD. (c) After label and conditional shift correction, i.e., the GLS correction, predictor trained on source domain can be generalized to target domain. Better view in color.
@@ 3.3 Label Probability Distribution Estimation
$=\frac{\mathbf{1}_{a\times b}(y)~{}f_{\mathrm{Gau}}(y;\mu,{\Sigma})}{F_{\mathrm{Gau}}(v_{22};\mu,{\Sigma})-F_{\mathrm{Gau}}(v_{12};\mu,{\Sigma})-F_{\mathrm{Gau}}(v_{21};\mu,{\Sigma})+F_{\mathrm{Gau}}(v_{11};\mu,{\Sigma})}$
where $y=(y_{1},y_{2})^{T}$ is the gaze value, $\mathbf{1}_{a\times b}(y)$ is the characteristic function. Let $v_{11}=(a_{1},b_{1})^{T}$,$v_{12}=(a_{1},b_{2})^{T}$,$v_{21}=(a_{2},b_{1})^{T}$,$v_{22}=(a_{2},b_{2})^{T}$ denote the vertices of the rectangular range $(a_{1},a_{2})\times(b_{1},b_{2})$, $a\coloneqq(a_{1},a_{2})$ and $b\coloneqq(b_{1},b_{2})$ denote the intervals, $f_{\mathrm{Gau}}(y;\mu,{\Sigma})$ and $F_{\mathrm{Gau}}(y;\mu,{\Sigma})$ denote the probability and cumulative density functions of bivariate Gaussian distribution with mean $\mu$ and covariance $\Sigma$. By estimating the statistics $(\mu,\Sigma)$ with target pseudo label $\hat{y^{t}}=h(g(x^{t}))$, the importance weight function takes the form
@@ 3.4 Probability-Aware Conditional Alignment
Such change in estimation differs the distribution embedding. As we aim to align $P^{t}_{Z|Y}$ with the reweighted $P^{s}_{Z|Y^{\omega}}$ rather than $P^{s}_{Z|Y}$, it is necessary to derive a probability-aware estimation of COD, denoted by PCOD. We carefully complete the derivation, which involves extensive probability and matrix analysis. Here we directly present the results of the derivation and provide the details of the derivation process in Appendix B.
@@ 4 Experiment
Setups. We conduct experiments on standard CDGE datasets: ETH-XGaze ($\mathcal{D}_{E}$zhang2020eth ,Gaze360 ($\mathcal{D}_{G}$kellnhofer2019gaze360 ,MPIIFaceGaze ($\mathcal{D}_{M}$zhang2017mpiigaze and EyeDiap ($\mathcal{D}_{D}$funes2014eyediap . The face image data are normalized according to the preparation process summarized in cheng2024appearance . Details about these datasets are provided in Appendix A.1. Following previous works, we adopt $\mathcal{D}_{E}$, $\mathcal{D}_{G}$ as source domains and $\mathcal{D}_{M}$, $\mathcal{D}_{D}$ as target domains, as the former two domains have wider gaze distributions than the latter. Then four CDGE tasks are established: $\mathcal{D}_{E}\to\mathcal{D}_{M}$, $\mathcal{D}_{E}\to\mathcal{D}_{D}$, $\mathcal{D}_{G}\to\mathcal{D}_{M}$ and $\mathcal{D}_{G}\to\mathcal{D}_{D}$. Implementation details are provided in Appendix A.3. Each evaluation experiment is repeated five times and the average results are reported.
@@ 4.2 Applicability Analysis
Figure 4: Visualization of GLS correction process. For scatter plot (a)(c)(d)(f), label variables are all denoted by $\bullet$’ regardless of value and are distinguished from other domains by colors. In contrast, for t-SNE figure (b) and (e), label values are denoted by color gradients and source domain features are denoted by $\bullet$’ while target features are denoted by ’$+$
@@ A.2 Algorithm
1 Input: Source data $\{(\mathbf{x}_{i}^{s},\mathbf{y}_{i}^{s})\}_{i=1}^{n}$ and unlabeled target data $\{\mathbf{x}_{i}^{t}\}_{i=1}^{n}$, source pretrained model $(g,h)$;
@@
A.3 Implementation details
For experiments using ResNet-18 and ResNet-50 as backbones models, the ResNet model pre-trained on ImageNet is assembled with a two-layers MLP as shallow feature extractor and a linear layer as the gaze predictor. The deep backbone module of ResNet is frozen in cross-domain training process. We use the Adam optimizer with the learning rate of $53e-$ 5 and a cosine annealing scheduler to decrease the learning rate in the training process. The batch size is set to be 100. As the domain shift is distinct at the beginning, we alternately correct the label shift and the conditional shift to produce better pseudo label. The confidence that decides the truncated area in label shift correction process is emprically set to 0.7 for all tasks.

### v2

@@ Abstract
Aiming to generalize the well-trained gaze estimation model to new target domains, Cross-domain Gaze Estimation (CDGE) is developed for real-world application scenarios. Existing CDGE methods typically extract the domain-invariant features to mitigate domain shift in feature space, which is proved insufficient by Generalized Label Shift (GLS) theory. In this paper, we introduce a novel GLS perspective to CDGE and modelize the cross-domain problem by label and conditional shift problem. A GLS correction framework is presented and a feasible realization is proposed, in which an importance reweighting strategy based on truncated Gaussian distribution is introduced to overcome the continuity challenges in label shift correction. To embed the reweighted source distribution to conditional invariant learning, we further derive a probability-aware estimation of conditional operator discrepancy. Extensive experiments on standard CDGE tasks with different backbone models validate the superior generalization capability across domain and applicability on various models of proposed method.
@@ 1 Introduction
Gaze estimation (GE) is crutial for understanding human attention in many areas, such as human-robot interaction yang2023natural , virtual reality wagner2024gaze and medical analysis ma2024eye . Recently, leveraging the capabilities of deep models, appearance-based gaze estimation zhang2015appearance ; nagpure2023searching ; liu2024test has attracted wide attention due to its low device requirements and end-to-end workflows. However, the great performance of deep models relies on the identical assumption of training and testing data distributions, which is hard to fulfill in real-world applications. When being applied across data domain, models’ performance usually degrade dramatically due to the domain shift caused by the changes of subjects, environment, etc. Therefore, dealing with the cross-domain gaze estimation (CDGE) problem is crutial for expanding application prospects of gaze estimation.
@@
Figure 1: Illustration of GLS correction framework. The DG and UDA methods typically learn the invariant representation across domains. Differently, the proposed GLS correction framework consider both the label shift correction and the probability-aware invariant representation learning. Note that the color of the "Alignment" component is split into orange and black to indicate that both GLS and UDA perform distribution alignment.
@@ 2 Preliminary
2.1 Problem Setup
Let $X$ and $Y$ be image and gaze variables defined on $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{D}^{s}=\{(x^{s},y^{s})\}$ and $\mathcal{D}^{t}=\{(x^{t},y^{t})\}$ be source and target domains with elements sampled from variables following different distributions, i.e., $(x^{s},y^{s})\sim(X^{s},Y^{s})\sim P^{s}_{XY}$ and $(x^{t},y^{t})\sim(X^{t},Y^{t})\sim P^{t}_{XY}$. For distribution $P$, the lowercase $p$ is the probability density function. Denote $g:X\to Z\in\mathcal{Z}$ as the feature transformation, $h:Z\to Y$ as the gaze predictor, then a learning model can be regarded as a tuple $(g,h)$. Let $\ell(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a loss function, then the CDGE problems aim to learn the model that minimizes the prediction error in target domain as
@@ 2.1 Problem Setup
2.2 Cross-domain Gaze Estimation (CDGE)
Existing CDGE methods can be roughly categorized as DG and UDA methods:
DG methods. Puregaze cheng2022puregaze and GazeCon xu2023learning tackle the CDGE problems by gaze feature purification bao2024unsupervised , which eliminate the gaze-irrelevant factors to obtain gaze-specific feature. GFAL xu2024gaze utilizes the gaze frontalization process and CUDA-GHR jindal2023cuda adopts gaze redirection yu2020unsupervised as auxiliary learning task to improve the generalization capability of features. An attractive advantage of DG methods is that they don’t need target domain samples and the gaze features can be generalized to unseen target domain. However, the improvement is usually limited due to the absence of knowledge from target domain.
UDA methods. Given some unlabeled samples from target domain, the UDA methods enhance the model performance in target domain by distribution alignment. DAGEN guo2020domain , PnP-GA liu2021generalizing and PnP-GA+ liu2024pnp align the source and target distributions by different metrics while CRGA wang2022contrastive and HFC liu2024gaze minimize the discrepancy in a contrastive learning paradigm. Recently, UnReGA cai2023source , DUCA zhang2024domain propose to extract domain-invariant representations by reducing the uncertainty within features. Although these DG and UDA methods differ in technology, they essentially focus on domain-invariant representation learning.
2.3 Generalized Label Shift (GLS) Theory
A joint distribution $P_{XY}$ can be factorized as $P_{XY}=P_{X|Y}P_{Y}$. Then the discrepancy between $P^{s}_{XY}$ and $P^{t}_{XY}$ is naturally deduced to the following two shifts:
$P^{s}_{Y}\neq P^{t}_{Y},$ (label shift) (2)
$P^{s}_{X|Y}\neq P^{t}_{X|Y}.$ (conditional shift) (3)
@@ 3 GLSGE: Perspective and Methodology
Here we provide a brief overview of the following sections. In Sec. 3.1, we introduce a novel GLS perspective to CDGE problem. In Sec. 3.2, we present the GLS correction framework and point out that existing DG and UDA methods can be viewed as partial realizations of this framework. In Sec. 3.3 and Sec. 3.4, we provide a feasible realization of the framework, in which a continuous importance reweighting strategy and a probability-aware estimation of COD are proposed to overcome the challenges arise from the continuity of gaze variable. In Sec. 3.5j, we summarize the proposed GLS correction method.
@@ 3.1 A GLS Perspective of CDGE problem
Figure 2: A GLS perspective of CDGE problem. (a) The label probability functions differ in functions support (colored area) and probability values (color degree), resulting the label shift between domains. (b) Conditional distributions of the same gaze label may differ due to factors like identity, background and illuminations.
@@
Conditional shift. The difference in sample level is more intuitive. In different scenarios the appearance image may differ significantly due to the background, illuminations, etc, leading to the discrepancy of conditional distributions, i.e.,
@@
Remark 3.1.
From a probabilistic perspective, both factorizations of the joint distribution—i.e., $P(X,Y)=P(X|Y)P(Y)$ or $P(Y|X)P(X)$ —are mathematically valid. However, we adopt the factorization $P(X,Y)=P(X|Y)P(Y)$ because it provides a more intuitive and interpretable modeling perspective for the CDGE task. Further discussion can be found in Appendix B.
@@ 3.2 GLS correction framework for CDGE
  • Estimating the importance weight function $\omega(y)$ and reweighted $P^{s}_{Y^{\omega}}\triangleq\omega(y)P^{s}_{Y}$ by
    $\omega=\underset{\omega}{\arg\min}\ \mathcal{L}_{\mathrm{lab}}(P^{s}_{Y^{\omega}},P^{t}_{Y}).$ (7)
@@
    $g=\underset{g}{\arg\min}\ \mathcal{L}_{\mathrm{cond}}(P^{s}_{Z|Y^{\omega}},P^{t}_{Z|Y}).$ (8)
@@
The terms $\mathcal{L}_{\{\}}(\cdot,\cdot)$ denotes the learning Objectives in each step. By considering different steps and distance on distributions, this framework is connected with existing CDGE methods.
@@
Figure 3: Overview of the proposed GLS correction method. (a) Reweighting source label distribution with bivariate Gaussian distribution estimated by pseudo target labels. The reweighted label distribution is utilized in both conditional alignment and task-specific training. (b) Conditional invariant representation learning. Two sets of conditional distribution are map to RKHS and the discrepancy is measured by the PCOD. (c) After label and conditional shift correction, i.e., the GLS correction, predictor trained on source domain can be generalized to target domain. Note that the text color reflects the component each learning objective is associated with.
@@ 3.3 Label Probability Distribution Estimation
$=\frac{\mathbf{1}_{a\times b}(y)~f_{\mathrm{Gau}}(y;\mu,{\Sigma})}{F_{\mathrm{Gau}}(v_{22};\mu,{\Sigma})-F_{\mathrm{Gau}}(v_{12};\mu,{\Sigma})-F_{\mathrm{Gau}}(v_{21};\mu,{\Sigma})+F_{\mathrm{Gau}}(v_{11};\mu,{\Sigma})},$
where $y=(y_{1},y_{2})^{T}$ is the gaze value, $\mathbf{1}_{a\times b}(y)$ is the characteristic function. Let $v_{11}=(a_{1},b_{1})^{T}$$v_{12}=(a_{1},b_{2})^{T}$$v_{21}=(a_{2},b_{1})^{T}$$v_{22}=(a_{2},b_{2})^{T}$ denote the vertices of the rectangular range $(a_{1},a_{2})\times(b_{1},b_{2})$, $a\coloneqq(a_{1},a_{2})$ and $b\coloneqq(b_{1},b_{2})$ denote the intervals, $f_{\mathrm{Gau}}(y;\mu,{\Sigma})$ and $F_{\mathrm{Gau}}(y;\mu,{\Sigma})$ denote the probability and cumulative density functions of bivariate Gaussian distribution with mean $\mu$ and covariance $\Sigma$. By estimating the statistics $(\mu,\Sigma)$ with target pseudo label $\hat{y}^{t}=h(g(x^{t}))$, the importance weight function takes the form
@@
Note that the truncated Gaussian distribution is only an alternative to model the gaze distribution. Other continuous distributions with compact support, such as exponential family distributions, uniform distributions, or even shallow neural networks, can also be utilized. The choice of $p_{Y}$ can be adapted to the characteristics of the specific task or dataset, e.g., in driver gaze datasets where gaze points tend to concentrate around a few specific directions like forward, mirrors, a truncated Gaussian mixture model may provide a more accurate fit.
@@ 3.4 Probability-Aware Conditional Alignment
Such change in estimation differs the distribution embedding. As we aim to align $P^{t}_{Z|Y}$ with the reweighted $P^{s}_{Z|Y^{\omega}}$ rather than $P^{s}_{Z|Y}$, it is necessary to derive a probability-aware estimation of COD, denoted by PCOD. We carefully complete the derivation, which involves extensive probability and matrix analysis. Here we directly present the results of the derivation and provide the details of the derivation process in Appendix C.
@@
As the PCOD is derived for the continuity and label shift of gaze variable, it is applicable to other tasks with similar distributional properties, providing a more general-purpose solution within the GLS correction framework. Of course, beyond distributional aspects, gaze estimation has other unique properties, such as rotational equivariance. Integrating these domain-specific priors into the GLS correction method is an interesting direction for future work.
@@ 3.5 GLS Correction Model for CDGE
Note that the introduced GLS correction framework is not limited to gaze estimation and can be potentially extended to a broad range of tasks with similar distributional properties such as pose estimation. We would also like to emphasize that our contribution is not only in the formulation of a general framework, but also in the design of a feasible and effective implementation for CDGE. We believe this step is crucial when adapting the general GLS framework to different tasks, and it often requires task-specific insights and custom solutions.
@@ 4 Experiment
Setups. We conduct experiments on four standard CDGE datasets: ETH-XGaze ($\mathcal{D}_{E}$zhang2020eth ,Gaze360 ($\mathcal{D}_{G}$kellnhofer2019gaze360 ,MPIIFaceGaze ($\mathcal{D}_{M}$zhang2017mpiigaze and EyeDiap ($\mathcal{D}_{D}$funes2014eyediap . The face image data are normalized according to the preparation process summarized in cheng2024appearance . Details about these datasets are provided in Appendix A.1. Following previous works, we adopt $\mathcal{D}_{E}$, $\mathcal{D}_{G}$ as source domains and $\mathcal{D}_{M}$, $\mathcal{D}_{D}$ as target domains, as the former two domains have wider gaze distributions than the latter. Then four CDGE tasks are established: $\mathcal{D}_{E}\to\mathcal{D}_{M}$, $\mathcal{D}_{E}\to\mathcal{D}_{D}$, $\mathcal{D}_{G}\to\mathcal{D}_{M}$ and $\mathcal{D}_{G}\to\mathcal{D}_{D}$. During cross-domain learning, we use 10% of the unlabeled target domain images for training and another 10% for validation, with the remaining 80% used for testing. It means that 4500 images in $\mathcal{D}_{M}$ and 1667 images in $\mathcal{D}_{D}$ are used for training in each task. Further discussion about the influence of target training data size is provided in Appendix A.3, while implementation details are provided in Appendix A.4. Each evaluation experiment is repeated five times and the average results are reported.
@@ 4.3 Ablation Analysis
Figure 4: Visualization of GLS correction process. For scatter plot (a)(c)(d)(f), label variables are all denoted by $\bullet$’ regardless of value and are distinguished from other domains by colors. In contrast, for t-SNE figure (b) and (e), label values are denoted by color gradients and source domain features are denoted by $\bullet$’ while target features are denoted by ’++
@@ 5 Conclusion
Acknowledgements
This work is supported in part by National Key R&D Program of China (2024YFA1011900), National Natural Science Foundation of China (Grant No. 62376291), Guangdong Basic and Applied Basic Research Foundation (2023B1515020004), Science and Technology Program of Guangzhou (2024A04J6413), and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (24xkjc013).
@@ A.2 Algorithm
1Input: Source data $\{(\mathbf{x}_{i}^{s},\mathbf{y}_{i}^{s})\}_{i=1}^{n}$ and unlabeled target data $\{\mathbf{x}_{i}^{t}\}_{i=1}^{n}$, source pretrained model $(g,h)$;
@@
A.3 Influence of target training data size
As mentioned in Sec. 4, we used 4500 images in $\mathcal{D}_{M}$ and 1667 images in $\mathcal{D}_{D}$ for cross-domain learning in each task. Here we provide the number of target domain samples used by various DA methods in our comparisons:
Method DAGEN PnP-GA GSA-Gaze HFC DCUA PnP-GA+
Num. 500 <100 1000 100 100 <100
Table 7: Number of target domain samples used by various DA methods
While the amount of unlabeled target-domain data we use may be larger than that used in some DG or UDA methods, we note the following:
  • The unlabeled data is inexpensive and easy to collect in real-world scenarios (e.g., by simply capturing images with a camera), whereas labeled data requires careful calibration and annotation. Therefore, using slightly more unlabeled samples does not significantly increase practical costs. Besides, their effective contribution to training is much lower than that of labeled data, so small differences in sample size have limited influence on performance.
  • Many existing approaches implicitly increase the diversity and amount of target-domain data via aggressive data augmentation, style transfer, or synthetic sample generation. Note that DG methods do not rely on target domain samples, but often heavily utilize data augmentation or auxiliary tasks to simulate domain shifts implicitly. While such methods implicitly enhance the effective size of training data, our method does not employ such augmentation strategies.
To address possible concern that performance gain might come from using more data, we conducted an experiment using only 100 unlabeled target samples, and report results below:
Method $\mathcal{D}_{E}$ $\mathcal{D}_{E}$ $\mathcal{D}_{G}$ $\mathcal{D}_{G}$ Avg
ResNet-18 8.05 9.03 7.41 8.83 8.33
GLSGE w/ $n^{t}>1000$ 5.31 6.21 5.43 7.30 6.06
GLSGE w/ $n^{t}=100$ 5.47 6.38 5.65 7.41 6.23
PnP-GA+ 5.34 5.73 6.10 7.62 6.20
Table 8: Comparision using only 100 unlabeled target samples with ResNet-18 as backbone model.
Although performance degrades slightly with fewer samples, our method still achieves SOTA performance. It’s also important to note that PnP-GA+, while using fewer samples, relies heavily on extensive data augmentations and up to 10 auxiliary models. In contrast, our method uses no data augmentation or auxiliary modules, making our results more interpretable and directly attributable to the core methodology.
Finally, we briefly explain why our method remains effective even with fewer target samples. It’s note that unlabeled data provides distributional information rather than supervision. In our method, label shift correction does not involve learning process and requires only a moderate number of samples for direct estimation. Conditional shift correction relies on feature alignment, which is more sample-sensitive and accounts for most of the performance drop. Nonetheless, even with a limited number of samples, our method achieves competitive results.
A.4 Implementation details
For experiments using ResNet-18 and ResNet-50 as backbones models, the ResNet model pre-trained on ImageNet is assembled with a two-layers MLP as shallow feature extractor and a linear layer as the gaze predictor. The deep backbone module of ResNet is frozen in cross-domain training process. We use the Adam optimizer with the learning rate of 3e53e-5 and a cosine annealing scheduler to decrease the learning rate in the training process. The batch size is set to be 100. As the domain shift is distinct at the beginning, we alternately correct the label shift and the conditional shift to produce better pseudo label. The confidence that decides the truncated area in label shift correction process is emprically set to 0.7 for all tasks.